Encircle the odd one of the following 
 (a) (–1) × (–1) (b) (–1) × (–1) × (–1) (c) (–1) × (–1) × (–1) × (–1) (d) (–1) × (–1) × (–1) × (–1) × (–1) × (–1) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (b) is correct here
Since,
(–1) × (–1) = 1
(–1) × (–1) × (–1) × (–1) = 1
(–1) × (–1) × (–1) × (–1) × (–1) × (–1) = 1 But, (–1) × (–1) × (–1) = -1
Hence, (–1) × (–1) × (–1) is the odd one.

Thanks